Assume that a person takes 0.100 g antacid. How many mol MgCl2 will be produced from ingested antacid, assuming all of the antacid undergoes the reaction? (M of Mg(OH)2 = 58 g/mol)


1
Expert's answer
2021-03-22T06:29:35-0400

Mg(OH)2 is the antacid; m(Mg(OH)2) = 0.100 g;

Molar mass Mr(Mg(OH)2) = Ar(Mg) + 2Ar(O) + 2Ar(H) = 24.305 g/mol + (2×16 g/mol) + (2×1 g/mol) = 58.305 g/mol.

n(Mg(OH)2) = m(Mg(OH)2)/Mr(Mg(OH)2) = (0.100 g)/(58.305 g/mol) = 0.0017 mol;

n(MgCl2) = n(Mg(OH)2) (stoichiometric coefficients);

n(MgCl2) = 0.00172 mol